3. Here are simplified financial statements for Phone Corporation in a recent year: INCOME STATEMENT (Figures in millions of dollars) Net sales Cost of goods sold Other expenses 13,193 4,060 Earnings before interest and taxes (EBIT) 2,566 Interest expense Income before tax Taxes (at 35%) Net income 685 658 1223 856 BALANCE SHEET (Figures in millions of dollars) End of YearStart of Year Cash and marketable securities 2382 238 932 3525 3,818 19973 19,915 4216 3,770 27714 27,503 Other cument assets Total current asset Net property, plant, and equipment Other long-term assets Total asvets Liabilities and shaneholders equity 2564 3,040 Short-term debt Other cutrent iabilties 811 4,794 $,400 Total current labilities Long-term debt and leases Other long-term lablities Shareholders equity Total liabilities and shareholders equity 27714 27,503 Calculate the following financial ratios for Phone Corporation a. Return on equity (Use average equity) b. Days in inventory (Use beginning inventory) e Operating profit margin (Use after-tax operating income) d. Total debt ratio (Use end of year values) e Current ratio (Use end of year values)
Answer #1

(a): Return on equity = net income/average shareholder’s equity = 1223/(9724+9121)/2 = 1223/9422.5

=12.98%

(b): Days in inventory = beginning inventory/cost of goods sold*365

= 238/4060*365

= 21.40

(c): Operating profit margin = after tax operating income/net sales

After tax operating income = operating income*(1-tax rate)

Operating income = EBIT = 2566 and tax rate = 35%. Thus after tax operating income = 2566*(1-35%) = 1667.90

Thus operating profit margin = 1667.90/13193

= 12.64%

(d): Total debt ratio = Total liabilities (end figure)/Total assets (end figure)

= (27714-9724)/27714

= 0.6491 (0.65 rounded off to 2 decimal place)

(e): Current ratio = current assets/current liabilities

= 3525/4794

= 0.7353 (0.74 rounded off to 2 decimal places)